# Scattering polarimetry for correlative nerve fiber imaging and multi-modal analysis

### v2

@@ Abstract
Mapping the intricate network of nerve fibers is crucial for understanding brain function. Three-Dimensional Polarized Light Imaging (3D-PLI) and Computational Scattered Light Imaging (ComSLI) map dense nerve fibers in brain sections with micrometer resolution using visible light. 3D-PLI reconstructs 3D-fiber orientations, while ComSLI disentangles multiple directions per pixel. So far, these imaging techniques have been realized in separate setups. A combination within a single device would facilitate faster measurements, pixelwise mapping, cross-validation of fiber orientations, and leverage the advantages of each technique while mitigating their limitations. Here, we introduce the Scattering Polarimeter, a microscope that facilitates correlative large-area scans by integrating 3D-PLI and ComSLI measurements into a single system. Based on a Mueller polarimeter, it incorporates variable retarders and a large-area light source for direct and oblique illumination, enabling combined 3D-PLI and ComSLI measurements. Applied to human and vervet monkey brain sections, the Scattering Polarimeter generates results comparable to state-of-the-art 3D-PLI and ComSLI setups and creates a multi-modal fiber direction map, integrating the robust fiber orientations obtained from 3D-PLI with fiber crossings from ComSLI. Furthermore, we discuss applications of the Scattering Polarimeter for unprecedented correlative and multi-modal brain imaging.
@@
neuroimaging, fiber architecture, brain structure, multi-modal imaging, polarization microscopy, light scattering
@@ Introduction
In the following, 3D-PLI and ComSLI results are compared to measurements with state-of-the-art setups. Finally, potential applications for multi-modal parameter analysis are demonstrated.
@@ Results
Correlative 3D-PLI and ComSLI measurements were performed with the Scattering Polarimeter (Fig. 1a) on several regions of two different samples: (1) Vervet (Fig. 1b) – 60 µm coronal section from a vervet monkey brain (male, 2.4 years old, no neurological diseases) that contains the corpus callosum, the corona radiata, the cingulum, and the fornix. (2) Human (Fig. 1c) – 50 µm sagittal section from a body donor (80 years old, female, no neurological diseases) that includes Broca’s region which is associated with language processing [18], containing areas of gray and white matter. The resulting parameter maps were compared to those obtained from existing, state-of-the-art 3D-PLI and ComSLI setups (see Methods). With these parameter maps, a multi-modal fiber direction map was generated by combining data from 3D-PLI and ComSLI. This map displays the most reliable fiber direction signal per pixel from the available modalities, based on a classification algorithm that determines the expected fiber architecture for each image pixel. The multi-modal fiber direction map exploits the robustness of 3D-PLI and the ability to detect fiber crossings of ComSLI, increases the total number of evaluated pixels and displays up to three fiber directions. A detailed description of the Scattering Polarimeter and the investigated brain sections is given in the Methods. A comprehensive overview of the fiber direction maps obtained with 3D-PLI and ComSLI is provided in Supplementary Fig. 1.
@@
Multi-Modal fiber direction map
A multi-modal fiber direction map combines information from multiple modalities to achieve the most reliable fiber direction per pixel. The reliability of signals from 3D-PLI or ComSLI largely depends on sample properties like the species, the age (i.e. time since embedding), the thickness of the sample [20], and the tissue preparation, but also on the local fiber architecture. The construction of a multi-modal fiber direction map was based on a previous classification of image pixels based on the local white matter fiber architecture, as described in the following:
@@
Due to the previously discussed suboptimal signal-to-noise ratio in ComSLI, the most obvious criterion for the identification of fiber crossings – whether one or multiple fiber directions are found in a pixel – is currently not sufficient. However, multi-modal parameter analysis can still find reliable directional information based on other optical parameters such as average scattering and peak prominence, despite the shortcomings of single individual parameters maps:
@@
Construction of a multi-modal fiber direction map
The construction of a multi-modal fiber direction map was based on the pixel classification for different fiber architectures and on the obtained fiber direction in 3D-PLI, ComSLI, and their difference (Fig. 4c). Fiber directions from 3D-PLI were selected over those from Mueller polarimetry due to the superior signal-to-noise ratio, particularly in cortical regions. The fiber directions for all investigated samples are shown in Supplementary Fig. 1. The multi-modal fiber direction map for the vervet corona radiata is shown in Fig. 4d.
Figure 4e illustrates the step-wise procedure used to construct the multi-modal fiber direction map.
The three direction maps of the multi-modal fiber direction map were assembled as follows:
Figure 4: Generation of the multi-modal fiber direction map (shown exemplary for the corona radiate region of the vervet monkey brain section). (a) White matter classification. Pixels classified as white matter are shown in blue, pixels classified as gray matter in white. (b) Fiber crossings classification. As a subset of white matter pixels, pixels identified as crossing fibers are shown in blue, other pixels in white. (c) Fiber direction map from 3D-PLI (middle) and ComSLI (bottom) measured with the Scattering Polarimeter; the absolute difference between the direction from 3D-PLI and the closest direction (if there is more than one) obtained with ComSLI is shown at the top. (d) Multi-modal fiber direction map. (e) Step-by-step construction of the multi-modal fiber direction map: The image pixels are filled based on the pixel classification with the directions $\phi_{\textrm{PLI}}$ from 3D-PLI and $\phi_{\textrm{SLI,1}}$, $\phi_{\textrm{SLI,2}}$, $\phi_{\textrm{SLI,3}}$ from ComSLI. The construction of each multi-modal direction map based on the best signal is displayed in three rows. Major changes between the steps are highlighted by dashed ellipses in the image. Corona radiata (cr), corpus callosum (cc), gray matter (gm).
@@
    Fiber directions in crossing regions are found more reliably with ComSLI because 3D-PLI cannot differentiate multiple directions and only detects the average signal. For pixels that were identified as part of crossing regions, the directions $\phi_{\textrm{SLI,1}}$, $\phi_{\textrm{SLI,2}}$, and $\phi_{\textrm{SLI,3}}$ were chosen for the respective multi-modal direction map, e.g. in the corona radiata.
@@
    For parallel fibers, the sinusoidal curve of the 3D-PLI signal is typically more stable than the line profiles of the ComSLI signal. Additionally, the angular discretization is lower for 3D-PLI than for ComSLI. For these pixels, $\phi_{\textrm{PLI}}$ was chosen if it was similar to $\phi_{\textrm{SLI,i}}$ with $i\in[1,2,3]$. In this step, changes in the second and third multi-modal direction map were small and limited to statistical fluctuations.
@@
The three new multi-modal direction maps were assembled into a single, multi-colored fiber direction map, where each image pixel is represented by $\times 2$ color-coded subpixels. If one pixel contains two or three fiber directions, the 2x2 subpixels show two or three colors, respectively. All pixels without any detected fiber direction are displayed in black. They are mostly part of the background area or gray matter.
@@
Comparison between single-mode and multi-modal fiber direction maps
@@
Figure 5 compares the single-mode fiber direction maps obtained from ComSLI measurements with the Scattering Polarimeter, and the multi-modal fiber direction maps obtained from a combination with the 3D-PLI fiber direction maps for the vervet brain sample. The single-mode and the multi-modal vector maps (vector length weighed with the corresponding average map and vectors overlaid for kernels of 20 $\times$ 20 pixels) are compared for three selected areas indicated in the fiber direction maps. The multi-modal fiber direction maps do not only provide information about multiple fiber directions obtained from ComSLI but also exhibit a better signal-to-noise ratio due the data from 3D-PLI, particularly for parallel fibers and gray matter, compared to the single-mode fiber direction maps obtained only with ComSLI. Noticeable improvements are observed in the gray matter, the corpus callosum, and parts of the cingulum. Additionally, background noise is inherently filtered out by the routine.
@@
Figure 5: Comparison between single-mode and multi-modal fiber direction maps for the vervet brain section. (a) Single-mode fiber direction maps obtained from a ComSLI measurement with the Scattering Polarimeter. Corona radiata (cr), corpus callosum (cc), cingulum (cg), fornix (fx), gray matter (gm), background (bg). Below, the corresponding multi-modal fiber direction maps obtained from a combination with the corresponding 3D-PLI measurement. Compared to the single-mode fiber direction map obtained only with ComSLI (first row), the multi-modal fiber direction map combines up to three fiber directions per pixel and exhibits a noticeably better signal-to-noise ratio. Background (bg) noise is inherently filtered out. (b) Single-mode vector maps for three selected regions indicated with 1, 2, 3 in a. The displayed vector length is weighed with the corresponding average scattering signal and vectors are visualized in kernels of 20 $\times$ 20 pixels. Below, the corresponding multi-modal vector maps with same visualization. The maps show more subtle differences between the single-mode directions obtained only with ComSLI and the multi-modal directions: Fiber directions are more distinct in the multi-modal vector map and less affected by noise that leads to additional fiber directions in 2 and 3 for parallel fibers in the single-mode vector map. Crossing fibers as in 1 do not change significantly because they are mostly retrieved from ComSLI alone.
@@
More subtle differences are evident in the vector maps: While the crossing regions in 1 do not differ significantly between the single-mode and the multi-modal fiber direction maps (as expected, as they are all retrieved from the ComSLI directions), the directions of the parallel in-plane fibers in the corpus callosum indicated by 2 are more distinct and less affected by noise. Although a second green direction manifests in the single-mode fiber direction maps, the multi-modal fiber direction maps show only the dominant magenta direction for most pixels. The parallel fibers of the corpus callosum area bordering the gray matter in 3 are improved even more visibly, nearly removing the additional yellow direction. Overall, the multi-modal fiber direction maps provide more detailed information with improved quality.
@@
However, a notable phenomenon occurs in the cingulum. Here, where the fiber bundles run nearly orthogonal to the sectioning plane, many pixels remain black in the multi-modal fiber direction maps, even when 3D-PLI clearly detects a dominant direction. This occurs because ComSLI fails to reliably detect a single or multiple directions. As a result, these pixels are not filled with 3D-PLI data because the 3D-PLI fiber directions are not considered reliable based on information from ComSLI. Previous studies discussed how the inclination obtained from 3D-PLI is not necessarily reliable for very steep fibers and how a combination of ComSLI and 3D-PLI can improve the results in this context [15].
@@ Discussion
The Scattering Polarimeter has successfully advanced from the prototypical stage to a fully operational multi-modal device. The novel device integrates oblique illumination for ComSLI and vertical illumination for 3D-PLI into a single, multi-modal system that has maintained the flexibility to realize illumination with different wavelengths or the use of other polarization states to perform full Mueller polarimetry. Software for measurement and evaluation were developed. The current study focused on enabling combined measurements of 3D-PLI and ComSLI in a single device and is a first proof-of-concept study. Future studies are needed to further optimize the measurement routine and enable high-throughput measurements of many tissue samples. The most urgent hardware upgrade is a brighter LED light source. Especially the ComSLI reference measurements indicated that measurements with a brighter light source lead to a better signal and less noise. The development of a custom light source with integrated polarization state generator (PSG) is optional but must be considered when scatterometry ComSLI [14] or compressed sensing ComSLI [25] become part of the routine measurement. For high-speed measurement applications, the integration of a custom scanning table is mandatory. In its current implementation, the 3D-PLI and ComSLI measurements with the Scattering Polarimeter take considerably more time than their respective reference measurements, partly because of more optical elements in the light path. Future studies should look into how the measurement time can be further reduced, e.g., by reducing the number of images (the LMP3D uses 9 instead of 18 images) and reducing the exposure time when implementing a brighter light source and a more sensitive camera. Even if measurements with the Scattering Polarimeter take longer than for state-of-the-art setups, there is no extra time needed for exchanging tissue samples or for performing image registration, which significantly reduces the total time needed for correlative measurements and analysis.
@@
Scattering polarimetry was demonstrated for samples with notably different optical properties. The performance was investigated using exemplary tissue samples with different scattering and transmission properties. Sources of statistical and systematic errors were identified and suggestions for further improvement were made. The human brain sample most closely resembles common measurement conditions: In 3D-PLI, samples are measured days after embedding. In ComSLI, it is beneficial to wait several weeks or even months until the transmittance has increased but it is uncommon to measure a sample that was embedded years earlier like the vervet brain sample. The spectrum of tissue samples revealed the strengths and limitations of each measurement modality, and has highlighted potential challenges that can arise with different tissue properties. A parameter study could now identify the optimal measurement parameters for routine scattering polarimetry of common tissue samples. The primary goal would be to determine the default exposure time and required number of repetitions for all modalities to obtain a good signal-to-noise ratio. Recently embedded tissue required relatively longer exposure times to yield a sufficient signal, especially for ComSLI. The optimal time for a multi-modal measurement after embedding should be investigated through a long-term study. In the following, the key findings for each modality are discussed.
@@
The field of multi-modal parameter analysis is extensive, with the Scattering Polarimeter generating more than 30 different maps (around 15 for Mueller polarimetry, 4 for 3D-PLI, and 12 from SLIX for ComSLI), covering a wide range of optical parameters, not even including the three-dimensional visualization of fiber orientations and vector maps. This work provides a proof-of-principle of how multi-modal parameter analysis serves the imaging of nerve fiber architecture. It is meant to showcase the potential of scattering polarimetry and multi-modal parameter analysis. Generally, multi-modal parameter analysis would not necessarily require a Scattering Polarimeter when good image registration of different modalities is available. However, correlative measurements with pixel-precise alignment of images are very beneficial for resolving structures of nerve cells that occur at this level of detail and allow for much faster multi-modal measurement and analysis.
@@
Accurate tissue and fiber classification is crucial for multi-modal parameter analysis. Automatic approaches such as histogram-based brightness analysis or machine learning-based methods should be explored to generalize the classification to all sorts of unknown samples. A routine to assemble a multi-modal fiber direction map was developed, based on the established fiber direction map from ComSLI. When all measurement techniques are optimized and tissue and fiber classification routines are enhanced, the multi-modal fiber direction map is expected to yield even more valuable results.
@@
In this study, the Scattering Polarimeter was introduced, enabling correlative measurements between 3D-PLI, ComSLI, and potentially Mueller polarimetry due to its design as a Mueller polarimeter based on liquid crystal variable retarders. This allows for pixelwise mapping and cross-validation of fiber orientations in brain samples, leveraging the strengths of each technique. The performance of the device was evaluated using two exemplary brain sections with different transmitting and scattering properties, one more suitable for 3D-PLI and one more suitable for ComSLI, thus highlighting both the strengths and challenges of each measurement modality. Error analysis based on Mueller matrix calculus gave further insight into the imaging properties of the device. Suggestions for further improvements were made, such as the implementation of a brighter light source. As a proof of concept, a multi-modal fiber direction map was created, combining the advantages of both techniques.
@@
A possible application for correlative measurements and multi-modal parameter analysis is the investigation of neurodegenerative diseases: When myelin degenerates – often in the course of neurodegenerative diseases [22, 10, 23, 24] – the consequent birefringence change leads to an apparent change of fiber directions: The existence of myelin leads to a strong negative birefringence. When myelin degenerates, only the remaining axon exhibits a weaker and positive birefringence. However, the fiber directions obtained by ComSLI are not influenced by this phenomenon because scattering is mostly unaffected by the effect [17]. Multi-modal analysis can support the investigation of neurodegenerative diseases by comparing ComSLI and 3D-PLI directions.

### v3

# Scattering polarimetry enables correlative nerve fiber imaging and multimodal analysis

@@ Abstract
Mapping the intricate network of nerve fibers is crucial for understanding brain function. Three-Dimensional Polarized Light Imaging (3D-PLI) and Computational Scattered Light Imaging (ComSLI) map dense nerve fibers in brain sections with micrometer resolution using visible light. 3D-PLI reconstructs 3D-fiber orientations, while ComSLI disentangles multiple directions per pixel. So far, these imaging techniques have been realized in separate setups. A combination within a single device would facilitate faster measurements, pixelwise mapping, cross-validation of fiber orientations, and leverage the advantages of each technique while mitigating their limitations. Here, we introduce the Scattering Polarimeter, a microscope that facilitates correlative large-area scans by integrating 3D-PLI and ComSLI measurements into a single system. Based on a Mueller polarimeter, it incorporates variable retarders and a large-area light source for direct and oblique illumination, enabling combined 3D-PLI and ComSLI measurements. Applied to human and vervet monkey brain sections, the Scattering Polarimeter generates results comparable to state-of-the-art 3D-PLI and ComSLI setups and creates a multimodal fiber direction map, integrating the robust fiber orientations obtained from 3D-PLI with fiber crossings from ComSLI. Furthermore, we discuss applications of the Scattering Polarimeter for unprecedented correlative and multimodal brain imaging.
@@
neuroimaging, fiber architecture, brain structure, multimodal imaging, polarization microscopy, light scattering
@@ Introduction
In the following, 3D-PLI and ComSLI results are compared to measurements with state-of-the-art setups. Finally, potential applications for multimodal parameter analysis are demonstrated.
@@ Results
Correlative 3D-PLI and ComSLI measurements were performed with the Scattering Polarimeter (Fig. 1a) on several regions of two different samples: (1) Vervet (Fig. 1b) – 60 µm coronal section from a vervet monkey brain (male, 2.4 years old, no neurological diseases) that contains the corpus callosum, the corona radiata, the cingulum, and the fornix. (2) Human (Fig. 1c) – 50 µm sagittal section from a body donor (80 years old, female, no neurological diseases) that includes Broca’s region which is associated with language processing [18], containing areas of gray and white matter. The resulting parameter maps were compared to those obtained from existing, state-of-the-art 3D-PLI and ComSLI setups (see Methods). With these parameter maps, a multimodal fiber direction map was generated by combining data from 3D-PLI and ComSLI. This map displays the most reliable fiber direction signal per pixel from the available modalities, based on a classification algorithm that determines the expected fiber architecture for each image pixel. The multimodal fiber direction map exploits the robustness of 3D-PLI and the ability to detect fiber crossings of ComSLI, increases the total number of evaluated pixels and displays up to three fiber directions. A detailed description of the Scattering Polarimeter and the investigated brain sections is given in the Methods. A comprehensive overview of the fiber direction maps obtained with 3D-PLI and ComSLI is provided in Supplementary Fig. 1.
@@
Multimodal fiber direction map
A multimodal fiber direction map combines information from multiple modalities to achieve the most reliable fiber direction per pixel. The reliability of signals from 3D-PLI or ComSLI largely depends on sample properties like the species, the age (i.e. time since embedding), the thickness of the sample [20], and the tissue preparation, but also on the local fiber architecture. The construction of a multimodal fiber direction map was based on a previous classification of image pixels based on the local white matter fiber architecture, as described in the following:
@@
Due to the previously discussed suboptimal signal-to-noise ratio in ComSLI, the most obvious criterion for the identification of fiber crossings – whether one or multiple fiber directions are found in a pixel – is currently not sufficient. However, multimodal parameter analysis can still find reliable directional information based on other optical parameters such as average scattering and peak prominence, despite the shortcomings of single individual parameters maps:
@@
Construction of a multimodal fiber direction map
The construction of a multimodal fiber direction map was based on the pixel classification for different fiber architectures and on the obtained fiber direction in 3D-PLI, ComSLI, and their difference (Fig. 4c). Fiber directions from 3D-PLI were selected over those from Mueller polarimetry due to the superior signal-to-noise ratio, particularly in cortical regions. The fiber directions for all investigated samples are shown in Supplementary Fig. 1. The multimodal fiber direction map for the vervet corona radiata is shown in Fig. 4d.
Figure 4e illustrates the step-wise procedure used to construct the multimodal fiber direction map.
The three direction maps of the multimodal fiber direction map were assembled as follows:
Figure 4: Generation of the multimodal fiber direction map (shown exemplary for the corona radiate region of the vervet monkey brain section). (a) White matter classification. Pixels classified as white matter are shown in blue, pixels classified as gray matter in white. (b) Fiber crossings classification. As a subset of white matter pixels, pixels identified as crossing fibers are shown in blue, other pixels in white. (c) Fiber direction map from 3D-PLI (middle) and ComSLI (bottom) measured with the Scattering Polarimeter; the absolute difference between the direction from 3D-PLI and the closest direction (if there is more than one) obtained with ComSLI is shown at the top. (d) Multimodal fiber direction map. (e) Step-by-step construction of the multimodal fiber direction map: The image pixels are filled based on the pixel classification with the directions $\phi_{\textrm{PLI}}$ from 3D-PLI and $\phi_{\textrm{SLI,1}}$, $\phi_{\textrm{SLI,2}}$, $\phi_{\textrm{SLI,3}}$ from ComSLI. The construction of each multimodal direction map based on the best signal is displayed in three rows. Major changes between the steps are highlighted by dashed ellipses in the image. Corona radiata (cr), corpus callosum (cc), gray matter (gm).
@@
    Fiber directions in crossing regions are found more reliably with ComSLI because 3D-PLI cannot differentiate multiple directions and only detects the average signal. For pixels that were identified as part of crossing regions, the directions $\phi_{\textrm{SLI,1}}$, $\phi_{\textrm{SLI,2}}$, and $\phi_{\textrm{SLI,3}}$ were chosen for the respective multimodal direction map, e.g. in the corona radiata.
@@
    For parallel fibers, the sinusoidal curve of the 3D-PLI signal is typically more stable than the line profiles of the ComSLI signal. Additionally, the angular discretization is lower for 3D-PLI than for ComSLI. For these pixels, $\phi_{\textrm{PLI}}$ was chosen if it was similar to $\phi_{\textrm{SLI,i}}$ with $i\in[1,2,3]$. In this step, changes in the second and third multimodal direction map were small and limited to statistical fluctuations.
@@
The three new multimodal direction maps were assembled into a single, multi-colored fiber direction map, where each image pixel is represented by $\times 2$ color-coded subpixels. If one pixel contains two or three fiber directions, the 2x2 subpixels show two or three colors, respectively. All pixels without any detected fiber direction are displayed in black. They are mostly part of the background area or gray matter.
@@
Comparison between single-mode and multimodal fiber direction maps
@@
Figure 5 compares the single-mode fiber direction maps obtained from ComSLI measurements with the Scattering Polarimeter, and the multimodal fiber direction maps obtained from a combination with the 3D-PLI fiber direction maps for the vervet brain sample. The single-mode and the multimodal vector maps (vector length weighed with the corresponding average map and vectors overlaid for kernels of 20 $\times$ 20 pixels) are compared for three selected areas indicated in the fiber direction maps. The multimodal fiber direction maps do not only provide information about multiple fiber directions obtained from ComSLI but also exhibit a better signal-to-noise ratio due the data from 3D-PLI, particularly for parallel fibers and gray matter, compared to the single-mode fiber direction maps obtained only with ComSLI. Noticeable improvements are observed in the gray matter, the corpus callosum, and parts of the cingulum. Additionally, background noise is inherently filtered out by the routine.
@@
Figure 5: Comparison between single-mode and multimodal fiber direction maps for the vervet brain section. (a) Single-mode fiber direction maps obtained from a ComSLI measurement with the Scattering Polarimeter. Corona radiata (cr), corpus callosum (cc), cingulum (cg), fornix (fx), gray matter (gm), background (bg). Below, the corresponding multimodal fiber direction maps obtained from a combination with the corresponding 3D-PLI measurement. Compared to the single-mode fiber direction map obtained only with ComSLI (first row), the multimodal fiber direction map combines up to three fiber directions per pixel and exhibits a noticeably better signal-to-noise ratio. Background (bg) noise is inherently filtered out. (b) Single-mode vector maps for three selected regions indicated with 1, 2, 3 in a. The displayed vector length is weighed with the corresponding average scattering signal and vectors are visualized in kernels of 20 $\times$ 20 pixels. Below, the corresponding multimodal vector maps with same visualization. The maps show more subtle differences between the single-mode directions obtained only with ComSLI and the multimodal directions: Fiber directions are more distinct in the multimodal vector map and less affected by noise that leads to additional fiber directions in 2 and 3 for parallel fibers in the single-mode vector map. Crossing fibers as in 1 do not change significantly because they are mostly retrieved from ComSLI alone.
@@
More subtle differences are evident in the vector maps: While the crossing regions in 1 do not differ significantly between the single-mode and the multimodal fiber direction maps (as expected, as they are all retrieved from the ComSLI directions), the directions of the parallel in-plane fibers in the corpus callosum indicated by 2 are more distinct and less affected by noise. Although a second green direction manifests in the single-mode fiber direction maps, the multimodal fiber direction maps show only the dominant magenta direction for most pixels. The parallel fibers of the corpus callosum area bordering the gray matter in 3 are improved even more visibly, nearly removing the additional yellow direction. Overall, the multimodal fiber direction maps provide more detailed information with improved quality.
@@
However, a notable phenomenon occurs in the cingulum. Here, where the fiber bundles run nearly orthogonal to the sectioning plane, many pixels remain black in the multimodal fiber direction maps, even when 3D-PLI clearly detects a dominant direction. This occurs because ComSLI fails to reliably detect a single or multiple directions. As a result, these pixels are not filled with 3D-PLI data because the 3D-PLI fiber directions are not considered reliable based on information from ComSLI. Previous studies discussed how the inclination obtained from 3D-PLI is not necessarily reliable for very steep fibers and how a combination of ComSLI and 3D-PLI can improve the results in this context [15].
@@ Discussion
The Scattering Polarimeter has successfully advanced from the prototypical stage to a fully operational multimodal device. The novel device integrates oblique illumination for ComSLI and vertical illumination for 3D-PLI into a single, multimodal system that has maintained the flexibility to realize illumination with different wavelengths or the use of other polarization states to perform full Mueller polarimetry. Software for measurement and evaluation were developed. The current study focused on enabling combined measurements of 3D-PLI and ComSLI in a single device and is a first proof-of-concept study. Future studies are needed to further optimize the measurement routine and enable high-throughput measurements of many tissue samples. The most urgent hardware upgrade is a brighter LED light source. Especially the ComSLI reference measurements indicated that measurements with a brighter light source lead to a better signal and less noise. The development of a custom light source with integrated polarization state generator (PSG) is optional but must be considered when scatterometry ComSLI [14] or compressed sensing ComSLI [25] become part of the routine measurement. For high-speed measurement applications, the integration of a custom scanning table is mandatory. In its current implementation, the 3D-PLI and ComSLI measurements with the Scattering Polarimeter take considerably more time than their respective reference measurements, partly because of more optical elements in the light path. Future studies should look into how the measurement time can be further reduced, e.g., by reducing the number of images (the LMP3D uses 9 instead of 18 images) and reducing the exposure time when implementing a brighter light source and a more sensitive camera. Even if measurements with the Scattering Polarimeter take longer than for state-of-the-art setups, there is no extra time needed for exchanging tissue samples or for performing image registration, which significantly reduces the total time needed for correlative measurements and analysis.
@@
Scattering polarimetry was demonstrated for samples with notably different optical properties. The performance was investigated using exemplary tissue samples with different scattering and transmission properties. Sources of statistical and systematic errors were identified and suggestions for further improvement were made. The human brain sample most closely resembles common measurement conditions: In 3D-PLI, samples are measured days after embedding. In ComSLI, it is beneficial to wait several weeks or even months until the transmittance has increased but it is uncommon to measure a sample that was embedded years earlier like the vervet brain sample. The spectrum of tissue samples revealed the strengths and limitations of each measurement modality, and has highlighted potential challenges that can arise with different tissue properties. A parameter study could now identify the optimal measurement parameters for routine scattering polarimetry of common tissue samples. The primary goal would be to determine the default exposure time and required number of repetitions for all modalities to obtain a good signal-to-noise ratio. Recently embedded tissue required relatively longer exposure times to yield a sufficient signal, especially for ComSLI. The optimal time for a multimodal measurement after embedding should be investigated through a long-term study. In the following, the key findings for each modality are discussed.
@@
The field of multimodal parameter analysis is extensive, with the Scattering Polarimeter generating more than 30 different maps (around 15 for Mueller polarimetry, 4 for 3D-PLI, and 12 from SLIX for ComSLI), covering a wide range of optical parameters, not even including the three-dimensional visualization of fiber orientations and vector maps. This work provides a proof-of-principle of how multimodal parameter analysis serves the imaging of nerve fiber architecture. It is meant to showcase the potential of scattering polarimetry and multimodal parameter analysis. Generally, multimodal parameter analysis would not necessarily require a Scattering Polarimeter when good image registration of different modalities is available. However, correlative measurements with pixel-precise alignment of images are very beneficial for resolving structures of nerve cells that occur at this level of detail and allow for much faster multimodal measurement and analysis.
@@
Accurate tissue and fiber classification is crucial for multimodal parameter analysis. Automatic approaches such as histogram-based brightness analysis or machine learning-based methods should be explored to generalize the classification to all sorts of unknown samples. A routine to assemble a multimodal fiber direction map was developed, based on the established fiber direction map from ComSLI. When all measurement techniques are optimized and tissue and fiber classification routines are enhanced, the multimodal fiber direction map is expected to yield even more valuable results.
@@
In this study, the Scattering Polarimeter was introduced, enabling correlative measurements between 3D-PLI, ComSLI, and potentially Mueller polarimetry due to its design as a Mueller polarimeter based on liquid crystal variable retarders. This allows for pixelwise mapping and cross-validation of fiber orientations in brain samples, leveraging the strengths of each technique. The performance of the device was evaluated using two exemplary brain sections with different transmitting and scattering properties, one more suitable for 3D-PLI and one more suitable for ComSLI, thus highlighting both the strengths and challenges of each measurement modality. Error analysis based on Mueller matrix calculus gave further insight into the imaging properties of the device. Suggestions for further improvements were made, such as the implementation of a brighter light source. As a proof of concept, a multimodal fiber direction map was created, combining the advantages of both techniques.
@@
A possible application for correlative measurements and multimodal parameter analysis is the investigation of neurodegenerative diseases: When myelin degenerates – often in the course of neurodegenerative diseases [22, 10, 23, 24] – the consequent birefringence change leads to an apparent change of fiber directions: The existence of myelin leads to a strong negative birefringence. When myelin degenerates, only the remaining axon exhibits a weaker and positive birefringence. However, the fiber directions obtained by ComSLI are not influenced by this phenomenon because scattering is mostly unaffected by the effect [17]. Multimodal analysis can support the investigation of neurodegenerative diseases by comparing ComSLI and 3D-PLI directions.